From Mon Jul  7 23:20:23 2003
Date: Thu, 27 Mar 2003 13:45:02 -0600 (CST)
Subject: SPARO update #3


This is the third general update on our efforts to get the SPARO
experiment installed and working.  The last one was in February.  In that
month, we left Paolo with an "almost working" system.  But after about
five weeks of hard work, he is now at the threshhold of having a fully
operational experiment.

There were two problems with the experiment at station closing.  The first
was that we never had time during summer to implement real-time pointing
correction for base tilt.  This work has now been finished, and we now
remove (in real time) a base tilt of approx. 0.9 arcminutes amplitude.  
We are now tracking sources with an accuracy of 0.5 arcmin.

We left Paolo with another problem, although we did not know about it at
station closing.  During the last few days of the season, we had installed
a camera to monitor the snow buildup on the condenser.  The observing
scripts were modified so as to cause a photo to be taken and saved to disk
every few minutes.  Unfortunately, there developed an intermittent problem
with this.  Specifically, the use of this "framegrabber" device in the
scripts would sometimes cause the SPARO data to be not written to disk.
Greg, Matt, and Paolo worked hard to solve this problem.  It has now gone
away, hopefully for good.

The tracking has been generally excellent and stable.  However, there was
one major exception.  Around March 15th, we had a relatively sudden change
in the pointing, where the dec pointing changed by about 5-6 arcmin.  The
RA pointing seems to have also changed, at the same time, but by a much
smaller amount.  The RA pointing change was so small that it is hard to be
sure that it really did change.  But it looks to me like it did change, by
about 1.5 arcmin (real arcmin).

This pointing change was a one-time event, and we have not experienced any
other pointing problems.  However, just after this pointing change, Paolo
noticed that the two chopper encoders were in disagreement by 200 mV.  
200 mV is 1.75 arcminutes.  we are not sure *when* the disagreement
between the two chopper encoders started.  it could have started any time
between mid-Feb and mid-March.  So the mid-March "pointing event" and the
encoder disagreement could be unrelated problems.  however, the near
coincidence between 1.5 arcmin and 1.75 arcmin suggests that they are the
same problem.

We are still in the process of diagnosing this "pointing event" problem.  
The possibilities we have thought of are:

(1) some change in position of chopper, e.g. loose bolts on chopper plate
or other part of chopper.  I hear that something like this happened to
Matt with similar effect last winter.

(2) slipped coupling or flakey/bad bit on dec encoder.  this would not
explain an RA pointing change.  Dave Chuss is pulling together some code
to examine the telemetry so we can use this as a tool to study this

(3) some change inside SPARO (e.g. broken Kevlar support).  i have
carefully studied the cryostat logs.  there has been no change in any of
the temperatures inside SPARO, so it seems like we can rule out this

(4) software.  according the SPARO data files, the coords sent to comsoft
did not change.  so if software is the culprit, it would seem that the
problem is in comsoft.

The full story on the pointing measurements with SPARO can be seen at 
Once we have made further progress in diagnosing this "pointing event" we
will be in contact with some of you for possible help and advice.

In the meantime, we have started polarimetry.  We have now done a total of
about 20-30 hours of polarimetry, starting about a week ago.  The reason
for the low duty cycle is that the weather was bad, so we were doing some
final measurements of the tilt, and also Paolo was testing his new
observing scripts that have new features like dead-zone recognition,
ability to alternate between polarimetry and photometry, etc.  But all
appears to be working well.  The magnitude of polarization of the peak of
RCW 108 is very small.  The weather is improving today with tau dropping
to values near 1.5.  Our first science target is RCW 57.  Just as the
satelite was going away this AM, Paolo was doing final pointing checks in
preparation for starting polarimetry of this source.

If there is something i haven't mentioned in this e-mail, that means it's
been working very nicely (e.g., heaters).